OpenStudy (anonymous):

will medal and fan Indicate the equation of the given line in standard form. The line that contains the point Q( 1, -2) and is parallel to the line whose equation is y - 4 = 2/3 (x - 3)

OpenStudy (haseeb96):

The slope of the above line is 2/3 So line are parallel so their slope will be same...m=2/3 Use this formula y-y1=m(x-x1) Just put the point (1,-2) in the above equation at the place of (x1,y1)and also put m in that formula and show me your work with answer.